### v1

# A Unified Charge-Dependent Modulation Model for AMS-02 Proton and Antiproton Fluxes during Solar Minimum

Abstract
We develop a unified charge-dependent solar modulation model by solving the three-dimensional Parker transport equation, incorporating a realistic wavy heliospheric current sheet to treat drift effects self-consistently. Using a local interstellar spectrum from GALPROP constrained by Voyager data, we fit the model to time-resolved proton and antiproton fluxes measured by the Alpha Magnetic Spectrometer - 02 (AMS-02) during the solar-quiet period (May 2011 – June 2022). To enable rapid parameter scans, we employ neural-network-based surrogate models to compute propagation and modulation matrices efficiently. The results demonstrate that the model simultaneously describes the observed proton and antiproton fluxes with physically reasonable parameters, providing a unified account of charge-dependent modulation.
pacs:
96.50.sb, 96.60.-j
I Introduction
Galactic cosmic rays (GCRs) detected at the top of the atmosphere (TOA) originate from sources such as supernova remnants (SNRs). After propagating through the galaxy, they undergo modulation by the heliospheric magnetic field (HMF) upon entering the solar system. The injection, propagation and modulation processes collectively shape the GCR spectrum observed at TOA. Consequently, the interdependence of these processes complicates the interpretation of cosmic ray flux, especially in the GeV energy range, obscuring potential anomalous signals, such as secondary particles from dark matter with masses in the tens of GeV range linInvestigatingDarkMatter2019 ; cuiPossibleDarkMatter2017 ; yuanPropagationCosmicRays2017 ; zhuExplainingGeVAntiproton2022 .
Two strategies effectively disentangle the modulation process from the injection and propagation processes: examining the local interstellar spectrum (LIS) and studying temporal variations in the cosmic-ray (CR) spectrum. The CR LIS, entirely free of heliospheric modulation, solely reflects injection and propagation effects. However, only the two Voyager spacecrafts have traversed beyond the heliosphere to date cummingsGALACTICCOSMICRAYS2016 ; stoneCosmicRayMeasurements2019 . Their CR observations, spanning MeV to sub-GeV energies, remain the only available LIS measurements. Alternatively, temporal variations in the TOA spectrum can separate modulation effects, as the HMF varies over the 11-year solar cycle while injection and propagation processes remain constant. The AMS-02 and the Payload for Antimatter Matter Exploration and Light-nuclei Astrophysics (PAMELA) experiments have provided precise, time-resolved CR flux measurements over a full solar cycle adrianiTimeDependenceProton2013 ; amscollaborationTemporalStructuresElectron2023 ; amscollaborationTemporalStructuresPositron2023 ; aguilarPeriodicitiesDailyProton2021 ; amscollaborationAntiprotonsElementaryParticles2025 . Combined with Voyager’s sub-GeV LIS data, these measurements can be employed to illuminate modulation effects.
Upon encountering the heliosphere, cosmic ray (CR) particles diffuse through the HMF. The structure of HMF induces charge-dependent drifts, while the solar wind drives convection and adiabatic energy losses. Particles of opposite charges drift in opposing directions, resulting in distinct diffusion paths and modulation strengths. Phenomenologically, the force field approximation (FFA) describes this effect by assigning distinct modulation potentials to each charge sign gleesonSolarModulationGalactic1968 ; cholisPredictiveAnalyticModel2016 ; kuhlenTimeDependentAMS02ElectronPositron2019 ; lipariSpectralShapesFluxes2019 ; silverTestingCosmicRayPropagation2024 ; koldobskiyValidationNeutronMonitor ; wangTimeDependentSolar2019 . Within this framework, particles sharing the same charge sign have been found to exhibit consistent modulation behavior zhaoUnifiedSolarModulation2025 . Although the FFA with charge-dependent modulation potential phenomenologically characterizes drifts, it neither elucidates the underlying drift mechanisms nor predicts modulation differences induced by drift.
Using recently published time-resolved fluxes of protons, antiprotons, electrons, and positrons measured by the AMS-02 collaboration from May 2011 to June 2022, we investigate drift effects more realistically amscollaborationAntiprotonsElementaryParticles2025 . We numerically solve the three-dimensional Parker transport equation for CR diffusion in the HMF via a time-backward stochastic method jiangNewScenarioSolar2023 . This approach, consistent with the SolarProp software kapplSOLARPROPChargesignDependent2016 , also incorporates a realistic wavy heliospheric current sheet (HCS) structure to model HCS drift, a component distinct from curvature and gradient drifts in the HMF peiThreedimensionalWavyHeliospheric2012 . By applying the LIS derived from the GALPROP package moskalenkoProductionPropagationCosmicRay1998 ; strongPropagationCosmicRayNucleons1998a , we demonstrate that this model provides a unified description of modulation effects for all the four elementary particle species, regardless of charge sign or mass, and is fitted to Voyager and AMS-02 measurements. Such a unified model can be further applied to study potential anomalous signals that are obscured in the GeV energy range of CR spectra.
This paper is structured as follows. Section III introduces the Parker transport equation, including the structure of the heliospheric magnetic field (HMF) and heliospheric current sheet (HCS), along with its stochastic solution. Section V presents the model parameter settings and fitting results. Section VI provides the discussion and conclusions.
II Local Interstellar Spectra
After being accelerated at their sources, CR particles propagate through the Galactic magnetic field, undergoing energy losses, fragmentation, convection, and reacceleration processes. These particles eventually reach the local interstellar region before encountering the Solar System. The LIS within this region serves as the fundamental input that determines the modulated spectra observed at the TOA.
The LIS has been measured by the two Voyager spacecraft directly, which remain the only human-made objects that have crossed the heliopause and exited the heliosphere. Their Cosmic Ray Subsystem (CRS) detects electrons/positrons in the energy range of 3-110 MeV and nuclei from 1-500 MeV/nucleon. However, as this system is unable to discriminate particle charge signs, Voyager data cannot determine the LIS of secondary species such as antiprotons and positrons, despite providing crucial direct measurements of primary cosmic rays like protons.
Direct measurements of local interstellar cosmic rays (CRs) are unavailable for energies above 0.5 GeV, necessitating a model to represent the local interstellar spectrum (LIS) in this range. For primary species such as protons and electrons, a spline interpolation model spanning MeV to tens of GeV can be applied to describe the LIS. In this work, however, we aim to develop a unified description for not only these primary particles but also secondary species, specifically antiprotons and positrons. To achieve a physically consistent model for all these particles, we utilize the GALPROP package.
In the GALPROP model, the injection of CRs into the Galaxy is described by a triple broken power-law spectrum, defined by spectral indices $\nu_{1}$, $\nu_{2}$, and $\nu_{3}$, along with break rigidities $R_{\mathrm{br,1}}$ and $R_{\mathrm{br,2}}$. After injection, CRs propagate through the Galactic halo of half-height $z_{h}$, undergoing diffusion and reacceleration before reaching the local interstellar medium. This reacceleration is characterized by the Alfvén speed $v_{A}$. The diffusion coefficient is governed by the particle speed $\beta$ (in natural units) and rigidity $R$, in relation
$D=D_{0}\beta^{\eta}\begin{cases}\left(\dfrac{R}{R_{0}}\right)^{\delta},&R<R_{h},\\ \left(\dfrac{R_{h}}{R_{0}}\right)^{\delta}\left(\dfrac{R}{R_{h}}\right)^{\delta_{h}},&R\geq R_{h},\end{cases}$ (1)
where the reference rigidity $R_{0}$ is fixed at 4GV4\ $\mathrm${GV}. The broken-power-law form is introduced primarily to accommodate the high-energy boron-to-carbon (B/C) ratio measurements from AMS-02 and DAMPE genoliniCosmicrayTransportAMS022019 ; maInterpretationsCosmicRay2023 . In the present analysis, however, we do not consider data in that high-energy regime. We therefore exclude B/C measurements above 200 GeV and, correspondingly, omit the high-energy break parameters $R_{h}$ and $\delta_{h}$ in the following analysis.
For this work, we adopt the ranges of these injection and propagation parameters from the recent study maInterpretationsCosmicRay2023 , which constrained them using DAMPE measurements. The corresponding parameter ranges are listed in Table 1.
Parameter Range Units
$\nu_{1}$ [1.5 , 2.5 ] -
$\nu_{2}$ [2.2 , 2.5 ] -
$\nu_{3}$ [1.5 , 3.0 ] -
$R_{\mathrm{br,1}}$ [1 , 80 ] GV
$R_{\mathrm{br,2}}$ [100 , 10000] GV
$D_{0}$ [1 , 20 ] $\mathrm{10^{28}cm^{2}\,s^{-1}}$
$\eta$ [-3 , 1.5 ] -
$\delta$ [0.3 , 0.6 ] -
$v_{A}$ [6 , 80 ] $\mathrm{km\,s^{-1}}$
$z_{h}$ [1.5 , 15 ] kpc
Table 1: Allowed ranges for Galactic CR injection and propagation parameters.
III Modulation Model
The CRs entering the heliosphere encounter the heliospheric magnetic field (HMF), which is embedded within the solar wind. Their transport is governed by Parker’s equation parkerPassageEnergeticCharged1965 :
$\frac{\partial f(\mathbf{r},\mathbf{p},t)}{\partial t}=$ $\nabla\cdot\left(\mathbf{K}\cdot\nabla f(\mathbf{r},\mathbf{p},t)-\mathbf{V_{SW}}f(\mathbf{r},\mathbf{p},t)\right)$ (2)
$+\frac{1}{3}(\nabla\cdot\mathbf{V_{SW}})\frac{\partial f(\mathbf{r},\mathbf{p},t)}{\partial\ln\,p}$
where $f(\mathbf{r},\mathbf{p},t)$ is the phase-space distribution function of galactic cosmic rays (GCRs). The three terms on the right-hand side of Equation (2) describe diffusion, adiabatic energy loss, and convection, respectively. The propagation of CRs in the HMF is governed by the tensor $\mathbf{K}$. The solar wind velocity, $\mathbf{V_{SW}}$, drives both the adiabatic energy loss and convection processes. We then describe the assumed solar wind velocity ($\mathbf{V_{SW}}$) distribution and corresponding HMF structure, along with the detailed diffusion model in the following subsections, to outline the implementation of our modulation model.
III.1 Heliosphere Structure
The solar wind, emanating from the Sun, extends magnetic field lines into interplanetary space, forming the HMF. Frozen into the solar wind plasma and rooted in the Sun’s northern and southern hemispheres, these field lines establish two distinct domains of opposite magnetic polarity—outward and inward, respectively. Their interface forms the heliospheric current sheet (HCS), a thin layer of electric current induced by the interaction of these oppositely directed magnetic fields. Due to the Sun’s rotation together with the tilt of its magnetic axis, both the HMF and HCS are drawn into spiral configurations: the field lines form the archetypal Parker’s spiral parkerDynamicsInterplanetaryGas1958 , while the HCS is warped into a complex, wavy structure resembling a ballerina’s skirt.
The large-scale structure of the spiral HMF can be described by
$\mathbf{B}=\frac{B_{0}r_{0}^{2}}{r^{2}}\left(\hat{\mathbf{e}}_{r}-\frac{\Omega_{\odot}r}{\mathrm{V_{SW}}}\sin\theta\hat{\mathbf{e}}_{\phi}\right)\left[1-2\mathrm{H}(\theta-\theta_{\mathrm{cs}})\right],$ (3)
where $B_{0}$ is the reference magnetic field strength at a distance $r_{0}$, $\Omega_{\odot}$ is the solar angular rotation velocity corresponding to a period of 27.5 days, and $\mathrm{H}(\theta-\theta_{\mathrm{cs}})$ is the Heaviside step function with $\theta_{\mathrm{cs}}$ representing the polar angle of the HCS. Gauss’s law for magnetism necessitates the conservation of the quantity $B_{0}r_{0}^{2}$ throughout the HMF. Transverse perturbations of the HMF near the Sun can significantly enhance the field magnitude in polar regions, modifying the archetypal spiral and introducing a $\hat{\mathbf{e}}_{\theta}$ component fichtnerCosmicRayModulation1996 ; jokipiiPolarHeliosphericMagnetic1989 . However, since only limited measurements of the polar HMF from Ulysses baloghHeliosphericMagneticField1995 are available and these modifications have minimal impact on modulation effects, we adopt the standard Parker spiral model in this work. Future observations from the second satellite of the Kuafu project tuIntroductionKuaFuProject2006 may warrant implementation of an updated HMF model.
The polar angle of the HCS $\theta_{\mathrm{cs}}$ varies with both radial distance $r$ and azimuthal angle $\phi$. This spatial dependence determines the HCS shape through the relation
$\tan\left(\frac{\pi}{2}-\theta_{\mathrm{cs}}\right)=\tan\alpha\sin\left[\phi+\frac{\Omega_{\odot}r}{\mathrm{V_{SW}}}-\Omega_{\odot}(t-t_{0})\right],$ (4)
where $\alpha$ is the tilt angle between the Sun’s magnetic and rotation axes. While this equation was originally approximated using a sine function for small tilt angles jokipiiEffectsDriftTransport1981 , we employ the correct tangent formulation following Ref. kotaEffectsDriftTransport1983 ; thomasEffectHeliosphericCurrent1986 ; peiThreedimensionalWavyHeliospheric2012 . Although $\alpha$ varies in time and space under realistic conditions, we adopt a constant value for simplicity in this work.
The solar wind velocity is radially aligned and has been observed to vary with heliocentric distance $r$ and colatitude $\theta$ during solar minimum bameUlyssesSolarWind1992 ; heberCosmicRaysHigh2006 . Inside the termination shock (TS), the wind speed $\mathrm{V_{SW}}$ remains constant with heliocentric distance, ranging from approximately $430\,\mathrm{km\,s^{-1}}$ at the equatorial plane to about $800\,\mathrm{km\,s^{-1}}$ in the polar regions. Beyond the TS, the wind speed decreases sharply, dropping to roughly $170\,\mathrm{km\,s^{-1}}$ at the equatorial plane. For the region extending from approximately $1\,\mathrm{AU}$ to the TS, we adopt the solar wind speed profile from Ref. potgieterModulationGalacticProtons2014 :
$\mathbf{V_{SW}}(r,\theta)=$ $V_{0}\left(1.475\mp 0.4\tanh\left[6.8\left(\theta-\frac{\pi}{2}\pm\theta_{0}\right)\right]\right)$ (5)
$\times\left(\frac{s+1}{2s}-\frac{s-1}{2s}\tanh\frac{r-r_{\mathrm{TS}}}{L}\right)\hat{\mathbf{e}}_{r},$
where $V_{0}=400\,\mathrm{km\,s^{-1}}$, $s=2.5$, $L=1.2\,\mathrm{AU}$, and the TS distance $r_{\mathrm{TS}}=90\,\mathrm{AU}$. The top and bottom signs correspond to the northern and southern hemispheres, respectively, with $\theta_{0}\equiv\alpha+15\pi/180$ defining the boundary between the slow and fast solar wind regions. Equation (5) yields wind speeds consistent with observations: approximately $430\,\mathrm{km\,s^{-1}}$ in the slow region near the equator and about $750\,\mathrm{km\,s^{-1}}$ in the fast region near the poles.
III.2 Diffusion Model
The propagation of CR particles through the HMF includes diffusion and drift. The diffusion, driven by the turbulent field component, is governed by the symmetric part of the diffusion tensor, $\mathbf{K}_{S}$. Conversely, the drift, resulting from the regular HMF component, can be described by the antisymmetric part, $\mathbf{K}_{A}$. Defined with respect to the direction of the HMF, the full diffusion tensor $\mathbf{K}$ is given by:
$\mathbf{K}$ $=\mathbf{K}_{S}+\mathbf{K}_{A}$ (6)
$=\begin{pmatrix}K_{\perp r}&&\\ &K_{\perp\theta}&\\ &&K_{\parallel}\end{pmatrix}+\begin{pmatrix}&-K_{A}&\\ K_{A}&&\\ &&0\end{pmatrix}$.
In the symmetric component of the diffusion tensor, $K_{\parallel}$ represents the diffusion coefficient parallel to the magnetic field, while $K_{\perp r}$ and $K_{\perp\theta}$ denote the perpendicular diffusion coefficients aligned with the radial and polar directions, respectively. This work adopts a typical empirical expression for $K_{\parallel}$ from Ref. potgieterModulationGalacticProtons2014 :
$K_{\parallel}=K_{0}\beta\left(\frac{B_{0}}{|\mathbf{B}|}\right)\left(\frac{R}{R_{0}}\right)^{a}\left[\frac{\left(\frac{R}{R_{0}}\right)^{m}+\left(\frac{R_{k}}{R_{0}}\right)^{m}}{1+\left(\frac{R_{k}}{R_{0}}\right)^{m}}\right]^{\frac{b-a}{m}}.$ (7)
Here, $K_{0}$ is a normalization factor of the order of $10^{23}~\mathrm{cm^{2}\,s^{-1}}$, $\beta\equiv v/c$ denotes the particle speed relative to the speed of light, $B_{0}$ is the reference magnetic field strength measured near Earth, and $R_{0}=1~\mathrm{GV}$ is the reference rigidity. The expression constitutes a smoothed broken power-law, characterized by power-law indices $a$ and $b$ below and above the break rigidity $R_{k}=3~\mathrm{GV}$, respectively, and a smoothing parameter $m=3$ that controls the sharpness of the transition.
The perpendicular diffusion has been realized to be anisotropic over the past two decades. The radial perpendicular diffusion coefficient is widely assumed to be proportional to the parallel diffusion coefficient via the relation $K_{\perp r}=0.02K_{\parallel}$. In contrast, numerical models constrained by the small latitudinal proton gradients observed by Ulysses during solar minimum indicate that the polar perpendicular diffusion coefficient is enhanced relative to the radial one heberCosmicRaysHigh2006 ; baloghGalacticAnomalousCosmic2008 . The polar perpendicular coefficient is commonly expressed as $K_{\perp\theta}=0.02K_{\parallel}f_{\perp\theta}$, where
$f_{\perp\theta}=A^{+}\mp A^{-}\tanh\left[-8(\left|90^{\circ}-\theta\right|\mp\theta_{F})\right],$ (8)
with $A^{\pm}=(d\pm 1)/2$, $\theta_{F}=35^{\circ}$, and an enhancement factor of $d=3$ towards the poles.
Defining the unit vector along the magnetic field as $\mathbf{e}^{B}$, the contribution from the antisymmetric component of the diffusion tensor can be reformulated as a cross product. Using index notation and noting that contraction with the symmetric part of the derivative vanishes, the expression simplifies as follows:
$\nabla\cdot(\mathbf{K}_{A}\cdot\nabla f)$ $=\partial_{i}\left(-\epsilon_{ijk}K_{A}\mathbf{e}^{B}_{k}\partial_{j}f\right)$ (9)
$=-\epsilon_{ijk}\partial_{i}(K_{A}\mathbf{e}^{B}_{k})\partial_{j}f-\epsilon_{ijk}K_{A}\mathbf{e}_{k}^{B}\partial_{i}\partial_{j}f$
$=\nabla\times(K_{A}\mathbf{e}^{B})\cdot\nabla f.$
This result indicates that antisymmetric diffusion produces an effect analogous to a drift velocity, defined by $\mathbf{V}_{d}\equiv\nabla\times(K_{A}\mathbf{e}^{B})$.
Under the weak-scattering assumption, the particle drift velocity is given as
$\mathbf{V}_{d}=\frac{\beta R}{3}\nabla\times\left(\frac{\mathbf{B}}{B^{2}}\right).$ (10)
Substituting the magnetic field from Equation (3) into Equation (10) yields the detailed expressions for the drift velocity in two distinct regions. For particles located away from the HCS, the drift velocity arises from the gradual gradient and curvature of the HMF jokipiiEffectsParticleDrifts1977 :
$\mathbf{V}_{d}^{\mathrm{GC}}=$ $\frac{2\beta Rr}{3B_{0}r_{0}^{2}(1+\gamma^{2})^{2}}\left[1-2\mathrm{H}(\theta-\theta_{\mathrm{cs}})\right]$ (11)
$\times\left(-\frac{\gamma}{\tan\theta}\hat{\mathbf{e}}_{r}+(2+\gamma^{2})\gamma\hat{\mathbf{e}}_{\theta}+\frac{\gamma^{2}}{\tan\theta}\hat{\mathbf{e}}_{\phi}\right),$
where $\gamma=r\Omega_{\odot}\sin\theta/\mathrm{V_{SW}}$. Near the HCS, the sharp magnetic field transition contributes an additional drift term. Assuming a locally flat current sheet, the magnitude of this term was derived and numerically computed burgerDriftTheoryCharged1985 . This numerical solution can be parameterized as burgerInclusionWavyNeutral1987 ; peiThreedimensionalWavyHeliospheric2012
$V_{d}^{\mathrm{HCS}}=$ $c\beta\mathrm{H}(2r_{g}-d)$ (12)
$\times\left[0.457-0.412\left(\frac{d}{r_{g}}\right)+0.0915\left(\frac{d}{r_{g}}\right)^{2}\right],$
where $d$ denotes the distance to the current sheet, $r_{g}$ is the Larmor radius, and the Heaviside function restricts this term to within twice the Larmor radius. Calculating the distance $d$ requires scanning the entire HCS. To achieve this efficiently, we employ an iterative one-dimensional optimization procedure in the $r$ and $\phi$ directions to locate the minimum distance. This algorithm typically converges in fewer than thirty iterations.
The HCS drift direction is perpendicular to the magnetic field and parallel to the current sheet. Considering the HCS shape described in Equation (4), the HCS drift component is given by:
$\mathbf{V}_{d}^{\mathrm{HCS}}=$ $V_{d}^{\mathrm{HCS}}\times\left[\frac{r\sin\theta_{\mathrm{cs}}\Omega_{\odot}}{V_{n}}\hat{\mathbf{e}}_{r}+\frac{\mathrm{V_{SW}}}{V_{n}}\hat{\mathbf{e}}_{\phi}\right$. (13)
$\left.-\frac{\tan\alpha\sin\theta_{\mathrm{cs}}\cos\phi_{0}(\mathrm{V^{2}_{SW}}+r^{2}\sin^{2}\theta_{\mathrm{cs}}\Omega_{\odot}^{2})}{\mathrm{V_{SW}}V_{n}}\hat{\mathbf{e}}_{\theta}\right],$
where $\phi_{0}=\phi+\Omega_{\odot}r/\mathrm{V_{SW}}-\Omega_{\odot}(t-t_{0})$, and $V_{n}$ is a normalization factor ensuring the vector inside the square bracket is a unit vector. Collectively, the gradient-curvature drift in Equation (11) and the HCS drift in Equation (13) describe all drift effects within the HMF.
Altogether, the modulation model comprises five parameters: the reference magnetic field $B_{0}$, the tilt angle $\alpha$, and three parameters ($K_{0}$, $a$, and $b$) that describe the diffusion coefficient. The parameter ranges, adopted from our previous work jiangNewScenarioSolar2023 , are provided in Table 2. The ranges for the spectral indices $a$ and $b$ have been updated to [0, 1] and [$a$, 3], respectively, to encompass their theoretically expected values derived from the power spectrum of the HMF jokipiiPropagationCosmicRays1971 .
Parameter Range Units
$B_{0}$ [3, 8] $\mathrm{nT}$
$\alpha$ [0, 75]
$K_{0}$ [0.001, 1.5] $\mathrm{10^{23}\,cm^{2}\,s^{-1}}$
$a$ [0, 1] -
$b$ [$a$, 3] -
Table 2: Allowed ranges for solar modulation parameters. These limits of $a$ and $b$ bracket the theoretical expectations of $a\approx 0.5$ and $b\approx 2$, derived from the magnetic-field power-spectrum analysis jokipiiPropagationCosmicRays1971 .
The three-dimensional Parker’s equation, with the complex structure described above, is solved using a backward-in-time stochastic differential equation (SDE) method. In this approach, pseudo-particles are traced backward from the Earth to the heliopause. The evolution of a particle’s position is governed by the SDE yamadaStochasticViewSolar1998 ; zhangMarkovStochasticProcess1999 ; koppStochasticDifferentialEquation2012 ; kapplSOLARPROPChargesignDependent2016 :
[…]
$\Delta p=-\frac{p}{3r^{2}}\frac{\partial r^{2}\mathrm{V_{SW}}}{\partial r}\Delta t.$ (15)
Tens of thousands of particles with varying energies were simulated using the SDE described above. The transition matrix of the spectrum was then computed statistically to characterize the modulation process. For this implementation, we have developed HELPROP [^1] https://github.com/linsujie/Helprop, a software package similar to SolarProp kapplSOLARPROPChargesignDependent2016 but specifically tailored to the present model.
The physically realistic HCS drift effect implemented in HELPROP plays a critical role in the particle transport within the heliosphere. As an illustration, Figure 1 shows the trajectories of two sample particles with rigidities near $1\,\mathrm{GV}$ propagating through the HMF during an A ¿ 0 magnetic polarity. Under this polarity, antiprotons experience a strong inward HCS drift, causing most antiprotons to enter the heliosphere primarily along the HCS. In contrast, protons are strongly drifted outward by the HCS drift, while the gradient‑curvature drift in the inner heliosphere tends to transport them toward the HCS. Consequently, only those protons that cross the polar regions without being diverted toward the HCS can avoid the strong outward HCS drift and ultimately reach Earth at 1 AU.
Figure 1: Representative particle trajectories simulated with HELPROP for A ¿ 0 polarity. The red and blue curves correspond to the proton and antiproton tracks, respectively. The gray wavy line denotes the HCS.
IV Model Fitting
This study aims to perform a global fit to determine the favoured parameter spaces of both the Galactic cosmic-ray propagation and solar modulation models. However, since a single GALPROP calculation requires several minutes and a HELPROP calculation can take tens of minutes, Monte Carlo sampling of the parameter space becomes computationally prohibitive. To address this challenge, we employ machine learning to train surrogate models that emulate these two processes, as detailed in Section IV.1.
Using these efficient emulators, we subsequently fit the AMS-02 and Voyager data. The configuration for this fitting procedure is described in Section IV.2.
IV.1 Machine Learning Model
The propagation of CRs in the Galaxy and their subsequent heliospheric modulation are fundamentally linear processes. This linearity implies that the influence of models such as GALPROP and HELPROP on the final CR spectrum can be effectively described by a linear transformation matrix, with each distinct set of model parameters corresponding to a unique matrix. To capture this relationship, we developed a dedicated artificial neural network (ANN) architecture capable of learning the mapping from parameters to matrix. This network is applied separately to each physical process, generating the precise transformation matrix for any given parameter set of the respective model, thereby explicitly preserving the linearity inherent in the propagation problem. We refer to this framework, which uses two separate networks to emulate GALPROP and HELPROP respectively, as ”PropMat”.
Figure 2: Deep learning architecture for spectral data processing in space physics, featuring propagation matrix (PropMat) generation and analysis across spatial domains (Galaxy and Heliosphere) using an encoder-decoder framework.
The architecture of PropMat is based on an encoder-decoder structure, as illustrated in Fig. 2. An initial Multi-Layer Perceptron (MLP) encoder maps the input model parameters into a latent representation, which extracts underlying features through nonlinear activation functions. This latent vector is then decoded to produce the final transformation matrix.
To handle multiple particle types, a subspace-based decoding scheme is implemented. The encoded latent vector is projected by a dedicated MLP into a particle-specific subspace of dimension $n_{\mathrm{sub}}$. The resulting vector serves as a set of coefficients, which are then fed into a coefficient-based matrix synthesis layer. In this layer, the coefficients linearly combine a set of learnable basis matrices to generate the final transformation matrix for the target particle type.
In this work, the subspace dimension $n_{\mathrm{sub}}$ is fixed at 32. This limited dimensionality constrains the matrix representation to a compact basis, encouraging the model to learn the principal features of the transformation. Such an approach is analogous to the Principal Component Analysis (PCA) algorithm jolliffePrincipalComponentAnalysis2014 , but substitutes its data-derived orthogonal basis with a set of learnable non-orthogonal basis matrices optimized for a custom objective. For physical transformation matrices, these principal features, typically limited in number, often correspond to low-frequency components. Consequently, the method inherently promotes smoothness in the learned transformations, ensuring a smooth output spectrum and thus supporting better generalization of the propagation result.
In order to train this ANN, we used the AdamWScheduleFree optimizer defazioRoadLessScheduled2024 for global optimization, followed by Stochastic Gradient Descent (SGD) bottouLargeScaleMachineLearning2010 for local fine-tuning. AdamWScheduleFree is a schedule-free variant of AdamW that eliminates the need for manual learning rate scheduling. Its interpolation-based update mechanism provides stable, competitive performance and enables rapid loss reduction during the initial training phase. In contrast, SGD is employed for the final fine-tuning stage. Its stochastic nature and simpler update rule help refine the model parameters and especially minimize prediction bias.
While the overall architecture of the two networks is similar, the loss functions for the GALPROP network and the HELPROP network are designed differently due to their distinct frameworks.
GALPROP employs discrete spatial and energy grids to model the evolution of particle densities during propagation. In this framework, the detailed transformation history of individual particles is not tracked. Consequently, GALPROP outputs only a final LIS spectrum without providing the corresponding transformation matrix. The artificial neural network (ANN) for GALPROP must therefore learn to infer this implicit transformation matrix, $M$, from a set of LIS spectrum samples. To enhance learning efficiency, we introduce a baseline matrix, $M_{\text{base}}$, obtained from multiple GALPROP runs with delta-function injections using a typical set of propagation parameters. Assuming the true transformation matrix shares a similar structure with $M_{\text{base}}$, we decompose the predicted matrix as
$M_{\text{pred}}^{ij}=\exp\Bigl(\log M_{\text{base}}^{ij}+\log M_{\text{dev}}^{ij}\Bigr).$ (16)
This allows the ANN to learn only the logarithmic deviation term, $M_{\text{dev}}$. To prevent the ill-conditioning of this term and mitigate over-fitting, we apply $L^{2}$ regularization specifically to $M_{\text{dev}}$. The corresponding loss function is therefore defined as the mean-squared error (MSE) between the predicted and true spectra, combined with this regularization term. Denoting the injection spectrum by $\psi_{I}$ and the output LIS spectrum by $\psi_{O}$, we have:
$\text{Loss}_{\text{G}}=\sum_{j}\Bigl(\log\frac{\sum_{i}M_{\text{pred}}^{ij}\,\psi_{\mathrm{I}}^{i}}{\psi_{\mathrm{O}}^{j}}\Bigr)^{2}+\sum_{ij}\beta\log\bigl(M_{\text{dev}}^{ij}\bigr)^{2},$ (17)
where the MSE is evaluated on a logarithmic scale because the spectral values span several orders of magnitude. The hyperparameter $\beta$, which controls the regularization strength, is assigned with a relatively large value at first and decreased gradually.
To generate the training data for this network, we employed a Markov Chain Monte Carlo (MCMC) algorithmforeman-mackeyEmceeMCMCHammer2013 within the parameter ranges specified in Table 1. The procedure was designed to sample propagation parameters that yield acceptable fits to AMS-02 boron-to-carbon (B/C) ratio amscollaborationPrecisionMeasurementBoron2016 , adopting HELPROP for solar modulation. Approximately 25,000 samples were generated, covering a broad region of parameter space where $\chi^{2}/\mathrm{d.o.f}<10$. These were split into a training set of 20,000 samples and a testing set of 5,000 samples. As shown in Figure 3, the trained model achieves a relative error between predicted and true spectra within 1% on the test set. This level of accuracy is below the measurement uncertainties of the AMS-02 cosmic-ray data and will not introduce significant errors in subsequent analyses.
Figure 3: Prediction accuracy of proton’s LIS, which is below the corresponding uncertainty of the AMS-02 experiment.
In contrast, HELPROP models heliospheric propagation by tracking a large ensemble of particles from Earth to the heliopause. Each simulation therefore intrinsically provides the corresponding transformation matrix $M_{\mathrm{O}}$. Although the matrix obtained from a finite particle sample is subject to statistical noise, we concentrate the sampling on a limited set of discrete energies to suppress fluctuations and then use interpolation to smoothly extend the matrix across the full energy range. Hence the ANN for HELPROP can learn the transformation matrix directly, leading to a straightforward loss function defined as the MSE between the predicted and target matrices:
$\mathrm{Loss}_{\mathrm{H}}=\sum_{ij}\Bigl(\log\frac{M_{\mathrm{pred}}^{ij}}{M_{\mathrm{O}}^{ij}}\Bigr)^{2},$ (18)
where, as in Equation (17), the MSE is also evaluated on a logarithmic scale. Since the typical energy change during heliospheric modulation is on the order of the modulation potential (approximately 1GeV1\ $\mathrm${GeV} for $|Z|=1$ particles), the energy-bin width in the transformation matrix must be considerably smaller than this scale. We therefore adopt a linear binning scheme with widths of 0.1GeV0.1\ $\mathrm${GeV} or 0.05GeV0.05\ $\mathrm${GeV}.
The training data for the HELPROP neural network were derived from the best-fit propagation parameters obtained during the GALPROP training phase. Using an MCMC approach with a likelihood function analogous to that used for GALPROP, we fitted jointly to the Voyager LIS proton flux and the monthly AMS-02 proton flux amscollaborationAntiprotonsElementaryParticles2025 by varying the injection and modulation parameters. Twenty Bartels rotations were selected from the monthly data. For each rotation, the magnetic field strength $B_{0}$ and the tilt angle $\alpha$ were fixed to their observed values (see Figure 5), and the MCMC sampled the parameter space $(K_{0},a,b)$. The individual sample sets from all twenty rotations were merged into a combined training dataset. In total, approximately 35,000 samples were generated and randomly split into a training set of 28,000 samples and a test set of 7,000 samples. On the test set, the trained model for HELPROP achieves a relative error between predicted and true spectra within 2% (see Figure 4). This accuracy is sufficient to avoid introducing significant computational errors in the following calculations.
Figure 4: Prediction accuracy of the modulated proton spectrum, which is below the corresponding uncertainty of the AMS-02 experiment.
Although the training samples correspond to discrete points in the $B_{0}$$\alpha$ plane, they are concentrated within a narrow band for the solar minimum period under study, thereby enabling the ANN to generalize effectively. We have verified this generalization capability by evaluating the model on months excluded from the training set, confirming its consistent performance throughout the entire period of interest.
Figure 5: Magnetic field strength $B_{0}$ and tilt angle $\alpha$ as functions of Bartels rotation. The curves represent measurements from Ref. smithACEMagneticFields1998 ; hoeksemaLargeScaleStructureHeliospheric1995 , with dashed segments corresponding to active periods and solid segments to solar minimum. Stars mark the rotations selected for training during solar minimum.
IV.2 Fitting Configuration
Utilizing the well-trained surrogate models, we conduct a global exploration of the parameter space using Markov Chain Monte Carlo (MCMC). The fit incorporates monthly proton and antiproton flux data from AMS, the AMS boron-to-carbon (B/C) ratio, and the local interstellar (LIS) proton spectrum from Voyager, bounded by the propagation and modulation parameter ranges given in Table 1 and Table 2.
To isolate charge-sign-dependent modulation effects, the analysis is confined to data from solar quiet periods, specifically May 2015 to June 2022, thereby avoiding the complications introduced during solar magnetic field reversal.
Each Bartels rotation involves three free modulation parameters, yielding a total of 276 parameters over the 92 rotations of interest. Performing a full Monte Carlo sampling in this high-dimensional space is highly inefficient. We therefore treat the modulation parameters as nuisance parameters. The MCMC sampling is performed only over the several injection and propagation parameters, while the modulation parameters for each Bartels rotation are optimized separately within each likelihood calculation. This reduces the effective sampling dimensionality to about ten, while still uniquely determining the best-fit modulation parameters for every rotation at each point in the reduced parameter space.
The observed CR fluxes at the TOA are shaped by source injection, Galactic propagation, and heliospheric modulation. With the Galactic propagation constrained by the B/C ratio, the Voyager LIS proton spectrum can effectively anchor the source injection spectrum for primary protons. The production of secondary antiprotons is then derived self-consistently from this framework. Given these constraints, simultaneously fitting a unified modulation model to the measured TOA proton and antiproton fluxes provides a direct test for the charge-sign-dependent drift effects inherent to our modulation model.
V Results
Figure 6: Best-fit proton (left) and antiproton (right) fluxes. The gray line represents the LIS, while the colored solid lines show the modulated spectra for the corresponding Bartels rotations. Colored dashed lines indicate the corresponding modulated spectra from the FFA, with the modulation potential $\phi$ given in the legend.
The fit to the AMS B/C ratio, the Voyager proton LIS, and the monthly AMS proton and antiproton fluxes yields a sensible fitting result. HELPROP, which incorporates a realistic HCS drift effect, successfully describes the modulation of both protons and antiprotons within a unified framework.
Figure 6 shows the best-fit local interstellar spectra for protons and antiprotons, along with the modulated spectra for several representative Bartels rotations, compared with Voyager and AMS-02 data. For comparison, the modulation result obtained using the FFA is plotted with dashed lines. The proton flux during Bartels rotation 2531 closely matches that during rotation 2543, whereas the antiproton flux is significantly higher in the latter. This difference results from the smaller tilt angle $\alpha$ in rotation 2543. During the studied period under positive magnetic polarity, as mentioned in Section III, protons reach Earth primarily through the polar region and are thus insensitive to changes in the tilt angle. In contrast, antiprotons detected at Earth are transported mainly via HCS drift. A larger tilt angle extends the drift path, enhances particle losses during transport, and consequently lowers the observed flux. Compared to the FFA, which requires distinct modulation potentials for the two Bartels rotations to account for such differences, our model incorporates tilt angle information and successfully reproduces the observed variations for both protons and antiprotons.
The best-fit propagation and injection parameters are listed in Table 3.
[…]
Table 3: Best-fit values, posterior medians (Med), and standard deviations ($\sigma$) for the injection and propagation parameters.
All propagation parameters are well-constrained by the B/C ratio together with a large set of antiproton and proton flux measurements. The normalization of the diffusion coefficient, $D_{0}=6.62\times 10^{28}\,\text{cm}^{2}\,\text{s}^{-1}$, and the size of the propagation halo, $z_{h}=7.21\,\text{kpc}$, are determined within broad yet acceptable ranges. The diffusion slope $\delta=0.40$ is close to the expectation from the Kolmogorov spectrum ($\delta=1/3$). The Alfvén speed $v_{A}=10.58\,\text{km}\,\text{s}^{-1}$ is consistent with typical interstellar medium values.
Only three of the five injection parameters are presented here. This is because the AMS-02 monthly data used in the fitting are measured at energies below 30 GeV, while the second injection break is expected to occur above several hundred GeV. Consequently, the parameters $R_{\mathrm{br,2}}$ and $\nu_{3}$ remain unconstrained by the data and were excluded from the fitting procedure. Figure 7 shows the one- and two-dimensional posterior distributions for the injection parameters $\nu_{1}$, $R_{\mathrm{br,1}}$, and $\nu_{2}$. The fit reveals a significant spectral break, with the index changing from $2.02$ to $2.21$ at a rigidity of approximately 2.45GV2.45\ $\mathrm${GV}. This break is effectively required by the spectral index of the Voyager LIS and the high-energy end of the AMS-02 data.
Figure 7: One- and two-dimensional posterior distributions for the injection parameters $\nu_{1}$, $R_{\mathrm{br,1}}$, and $\nu_{2}$, as constrained by the AMS monthly data and the Voyager LIS.
Figure 8 displays the time variation of the modulation parameter.
Figure 8: Time variation of the HELPROP parameters.
The reference diffusion coefficient at $1\,\mathrm{GV}$, $K_{0}$, always lies between $10^{22}\,\mathrm{cm^{2}\,s^{-1}}$ and $2\times 10^{22}\,\mathrm{cm^{2}\,s^{-1}}$, corresponding to a mean free path of several percent of an AU. This quantity is estimated over a wide range in different studies, using 2D, 3D, and semi-analytical models, from $\sim\!10^{22}\,\mathrm{cm^{2}\,s^{-1}}$tomassettiEvidenceTimeLag2017 ; songNumericalStudySolar2021 up to $\sim\!10^{23}\,\mathrm{cm^{2}\,s^{-1}}$potgieterModulationGalacticProtons2014 ; kuhlenTimeDependentAMS02ElectronPositron2019 . Our result falls within this expected range.
The low-energy power-law index $a$ of the diffusion coefficient ranges from 0.2 to 1, while the high-energy index $b$ ranges from 1 to 3. These ranges are consistent with expectations from power‑spectrum analyses of the HMF, which indicate that the maximum correlation length of HMF fluctuations is about 0.01 AU jokipiiPropagationCosmicRays1971 . Consequently, particles whose mean free path (MFP) is shorter than this correlation length experience an increase in diffusion coefficient similar to that in normal turbulence scenarios, such as the Kolmogorov or Bohm regimes, whereas particles with an MFP larger than this exhibit a much steeper rise in their diffusion coefficient.
The reduced $\chi^{2}$ for each Bartels rotation exhibits a significant correlation with the tilt angle $\alpha$. Periods with large $\alpha$ are associated with high solar activity and a complex magnetic field, which disturbs the modulation process. Consequently, the steady-state HELPROP model has difficulty to accurately describe these conditions, resulting in large reduced $\chi^{2}$ values. In extreme cases, such as around 2015 and 2022, these values can reach up to 10. Conversely, periods with a small $\alpha$ correspond to low solar activity, during which HELPROP provides a good description of the measurements. Adopting $\alpha<30^{\circ}$ as a criterion, the interval from 13 April 2017 to 19 September 2021 is identified as the solar quiet period. The total reduced $\chi^{2}$ over this interval is 0.58, indicating that HELPROP performs very well during the quiet period.
Both the HMF strength $B_{0}$ and the tilt angle $\alpha$ indicate a period of minimum solar activity around 2020. Such conditions, characterized by weak magnetic fields and low turbulence, result in weaker confinement of low-energy particles. This situation is reflected in the low-energy index $a$ of the diffusion coefficient, which reaches its minimum around 2020. Given that the normalization of the diffusion coefficient $K_{0}$ remains nearly constant, a smaller $a$ implies a larger parallel diffusion coefficient $K_{\parallel}$ in the sub-GeV energy range.
Thus, the solar minimum period is well described by HELPROP, and the behavior of the modulation parameter follows a reasonable trend within the model.
VI Conclusion
We have developed a unified charge-dependent solar modulation model to explain the monthly proton and antiproton fluxes measured by AMS-02 during solar quiet periods. In this model, the HELPROP, which incorporates a three-dimensional HCS drift effect, is employed to simulate drifts in a realistic manner, thereby accurately capturing the charge-dependent signatures in the modulation process.
Using the outputs of HELPROP and the galactic propagation code GALPROP, we separately constructed two surrogate models, called PropMat, based on a specially designed ANN architecture. Trained on tens of thousands of simulated samples, each surrogate model reproduces the corresponding energy-transformation matrix of GALPROP or HELPROP for any parameter set within the observationally allowed region, with a computational cost of only $\mathcal{O}(\mathrm{ms})$. The numerical error of these surrogate models is maintained at the one-percent level. They can therefore be used to perform efficient global parameter searches for various study objectives.
In this study, we employed the HELPROP surrogate model to rapidly fit the monthly AMS-02 proton and antiproton fluxes, thereby validating the modulation model. The LIS used in the modulation are derived from a GALPROP surrogate model and are further constrained by Voyager measurements. The fitting results demonstrate that HELPROP can accurately describe both the proton and antiproton monthly fluxes during the solar minimum period. Moreover, it naturally accounts for charge-dependent effects in a unified manner. The best-fit parameter values are physically reasonable, and their temporal evolution follows sensible trends.
This realistic drift-treatment approach should be applicable not only to protons and antiprotons, but also to electrons and positrons. In future work, we plan to apply the model to all these particle species using AMS-02 data during the solar minimum period. Furthermore, we will extend the model to periods with more complex heliospheric conditions.
This work is supported by the National Natural Science Foundation of China (NSFC) grants 12205388, 42150105, 12261141691, the Guangdong Provincial Key Laboratory of Advanced Particle Detection Technology (2024B1212010005), and the Guangdong Provincial Key Laboratory of Gamma-Gamma Collider and Its Comprehensive Applications (2024KSYS001). We thank the support from Fundamental Research Funds for the Central Universities, Sun Yat-sen University, No. 24qnpy123, and the Sun Yat-sen University Science Foundation.